OpenStudy (anonymous):

Complete the square for each expression. Write the resulting expression as a binomial squared. x^2 - 9x + ?

OpenStudy (anonymous):

Take 1/2 of the b term and square it In ax^2 + bx + c, the coefficient of x is the b term.

OpenStudy (anonymous):

so x^2 -9x=? x^2=-2*4 1/2*x+? ?=4 1/2^2*x - 81/4 (x-9/2)^2 ◀ right?